In a carnot cycle operating between 307 and 17 degree Celsius the maximum and minimum pressures are 60 2.4 and 1.04 bar.clclculate the cycle efficiency and the work ratio.assume air yo be working fluid.
1
Expert's answer
2020-01-16T07:38:42-0500

The cycle efficiency:


"\eta=1-\frac{T_2}{T_1}=1-\frac{290}{580}=0.5"

"WR=\frac{W_{net}}{W_{gross}}"

"W_{net}=(T_1-T_2)(S_1-S_2)"

4-a is isothermal, so


"S_A-S_4=R\ln{\frac{P_4}{P_2}}=0.287\ln{\frac{62.4}{1.04}}=1.175\frac{kJ}{kg}"

a-2 is isobaric, so


"S_2-S_a=c_p\ln{\frac{T_1}{T_2}}=1.005\ln{\frac{580}{290}}=0.6966\frac{kJ}{kg}"

"W_{net}=(580-290)(1.175-0.696)=139.2\frac{kJ}{kg}"

"W_{gross}=T_1(S_1-S_4)+c_v(T_1-T_2)"

"W_{gross}=580(0.48)+0.718(580-290)=486.62\frac{kJ}{kg}"

"WR=\frac{139.2}{486.62}=0.286"